# Giovanni Compiani Ilya Morozov Stephan Seiler

### v2

@@ 3.4.4 Summary
The Inmate (M) 0.325 Don’t Believe (S) 0.164 The Inmate (M) 0.163 The Housemaid (M) 0.170
The Housemaid (M) 0.250 Dopamine Detox (S) 0.149 The Housemaid (M) 0.149 The Inmate (M) 0.158
Don’t Believe (S) 0.088 The Inmate (M) 0.147 Don’t Believe (S) 0.145 Don’t Believe (S) 0.119
Art of Letting Go (S) 0.066 The Housemaid (M) 0.145 Dopamine Detox (S) 0.117 Serpent & Wings (F) 0.110
Serpent & Wings (F) 0.066 Art of Letting Go (S) 0.119 Art of Letting Go (S) 0.109 Art of Letting Go (S) 0.108
Dopamine Detox (S) 0.063 Serpent & Wings (F) 0.091 Court of Ravens (F) 0.099 Court of Ravens (F) 0.104
Court of Ravens (F) 0.058 Court of Ravens (F) 0.087 Serpent & Wings (F) 0.097 Dopamine Detox (S) 0.096
The Ritual (M) 0.056 The Ritual (M) 0.057 Ashes & Star (F) 0.070 Ashes & Star (F) 0.072
Ashes & Star (F) 0.029 Ashes & Star (F) 0.041 The Ritual (M) 0.050 The Ritual (M) 0.064
@@
Serpent & Wings (F) 0.275 The Housemaid (M) 0.155 Please Tell Me (M) 0.178 The Inmate (M) 0.184
Court of Ravens (F) 0.243 Don’t Believe (S) 0.155 The Inmate (M) 0.157 Please Tell Me (M) 0.166
The Inmate (M) 0.105 The Inmate (M) 0.153 The Housemaid (M) 0.134 The Housemaid (M) 0.123
The Ritual (M) 0.082 Please Tell Me (M) 0.139 Serpent & Wings (F) 0.116 Don’t Believe (S) 0.110
Please Tell Me (M) 0.080 Dopamine Detox (S) 0.116 The Ritual (M) 0.105 Court of Ravens (F) 0.108
The Housemaid (M) 0.070 Art of Letting Go (S) 0.100 Don’t Believe (S) 0.093 Dopamine Detox (S) 0.092
Don’t Believe (S) 0.052 Court of Ravens (F) 0.080 Dopamine Detox (S) 0.075 Serpent & Wings (F) 0.090
Art of Letting Go (S) 0.048 Serpent & Wings (F) 0.062 Art of Letting Go (S) 0.075 The Ritual (M) 0.065
Dopamine Detox (S) 0.045 The Ritual (M) 0.041 Court of Ravens (F) 0.066 Art of Letting Go (S) 0.063
@@
3.5 Implications for Product Substitution
So far, we have shown that our approach improves counterfactual predictions of second choices. One might wonder whether these $RMSE$ improvements are substantial and if they fundamentally change our understanding of product substitution in this market. To explore this further, in Table 1 we compare predicted second-choice probabilities with their counterparts observed in the experimental data.
Panel A examines substitution patterns for the self-help book Dopamine Detox. The data shows that the two other self-help books are, by far, its closest substitutes. The plain logit model misidentifies these substitutes, incorrectly predicting that people would switch to books Please Tell Me and The Inmate—two popular books with the largest market shares. The attribute-based mixed logit correctly identifies Don’t Believe Everything You Think as the closest substitute but mispredicts the second one, likely due to its over-reliance on estimated fixed effects. By contrast, our review-based model is the only one that correctly predicts all five closest substitutes in the correct order.
Panel B shows a similar example with the substitutes for the mystery book Please Tell Me. The plain logit model mispredicts second-choice probabilities, incorrectly suggesting that, in the absence of this book, participants would mainly switch to self-help books. By contrast, the attribute-based mixed logit model captures strong within-genre substitution, correctly identifying the top three closest substitutes. Our review-based model also correctly predicts the identities of these substitutes but swaps the first two. However, the review-based model recognizes that The Housemaid and The Inmate have significantly higher second-choice probabilities than the third-closest substitute—an insight the attribute-based model misses.
@@ 3.6 Implications for Pricing: Merger Simulations
Figure 5: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book, the figure shows the average price increase among the two merging books.
For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher, who sets prices by solving a joint first-order condition. In both cases, publishers take the prices of the other eight books as given, fixed at \$5.181818We fix the prices of the other books at \$5, the average value in the experimental dataset. We do not optimize prices of other books because the model does not include an outside option. Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles. However, our simulations offer a clear way to evaluate how estimated substitution patterns translate into pricing decisions. This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny. [^19] In 2022, a federal judge blocked the proposed merger between Penguin Random House and Simon & Schuster, citing concerns that it would stifle competition (U.S. Department of Justice, 2022). This merger would have reduced the “Big Five” publishers to the “Big Four” and significantly increased the concentration of the publishing market. Prior to that, Penguin Books and Random House merged in 2013 to form Penguin Random House—the largest trade book publisher globally.
@@
As discussed in Section 3.5, the plain logit model fails to capture strong within-genre substitution. Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones. This misclassification affects the predicted price changes in merger simulations. Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think. These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 9%—approximately 2 to 2.5 times higher.
The attribute-based mixed logit model produces predictions more aligned with the review-based model. In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases. However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to our selected text-based model.
@@ 4 Application to Online Retail Data
Next, we apply our approach to choice data from several online markets. We pursue two goals with this application. First, we aim to assess the external validity of our experimental results. Although our estimation approach works for books, this does not mean it would capture substitution patterns well in other product categories. By applying this approach across multiple categories, we also show that it can be applied broadly without being tailored to specific markets. Second, by exploring which text or image data best predict substitution patterns in various categories, we can offer practical guidance to empirical researchers on what unstructured data to collect for demand estimation.
@@ 4.1 Data
Online Purchases.
We use purchase data from the 2019-2020 Comscore Web Behavior Panel, which includes a sample of around two million U.S. households. We use the dataset constructed by Greminger et al. (2023), who classify over 12 million unique products from Amazon.com—browsed or purchased by Comscore panelists—into narrowly defined categories. This dataset also matches purchases with daily product price histories obtained from the third-party database Keepa.com. [^20] We do not observe product rankings, so we do not include them in our demand models. We focus on Amazon.com due to the high volume of Amazon purchases in the Comscore dataset.
We apply our method to 40 product categories. To construct this sample, we first select the 15 most-purchased products in each category and drop all other products. We then retain only the categories where these 15 products collectively account for at least 2,000 purchases. This selection criterion ensures that we observe a sufficient number of purchases to estimate both product-fixed effects and substitution patterns. Within each category, we treat each purchase as a choice instance and assume that consumers’ choice sets consist of the 15 selected products. [^21] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products.
The chosen product categories cover a diverse range of items such as clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation). The first column of Table 2 lists all included categories. On average, each category contains 662 purchases, with an average product price of \$43.
@@
Images and Texts.
We collect image and text data directly from product detail pages on Amazon.com. Figure 6 shows a sample product page. For image embeddings, we use the default product photo displayed at the top of the product detail page. For text embeddings, we gather information from several sections: product titles at the top, bullet points describing the product’s attributes (which we refer to as the “product description”), and the texts of the 100 most recent reviews for each product.
4.2 Estimation and Results
In each product category, we follow the same approach as in the experiment, extracting embeddings, reducing their dimensionality through PCA, and including them in the random coefficients logit model.
We do not observe product attributes in these categories. Collecting such data is challenging because each category has its own set of choice-relevant attributes. Further, we would need to subjectively decide which attributes to collect in each category—a challenge that originally motivated us to look at images and texts. For these reasons, we focus on comparing our estimates to those from the plain logit model. In other words, we focus on showing that texts and images contain information about product substitution and that our approach can reliably extract this information across many categories.
4.2.1 $AIC$ Improvements
Table 2 summarizes the estimation results, showing the selected model and data type for each category, along with the corresponding $AIC$ improvement relative to the plain logit model. Figure 7 further plots the distribution of $AIC$ improvements across all categories.
As a rule of thumb, we consider that a model has strong support compared to a simpler model if it lowers $AIC$ by at least $2.0$, and very strong support if it lowers $AIC$ by at least $5.0$. When a simpler model is nested within a more complex one with only one additional parameter, applying these $AIC$ thresholds is approximately equivalent to conducting a likelihood ratio test at the 5% and 1% significance levels. [^22] When nested models differ by one parameter, the reduction in $AIC$ is given by $\Delta AIC=2-\lambda_{LR}$, where $\lambda_{LR}$ is the likelihood ratio statistic. Therefore, our $\Delta AIC$ thresholds of $2.0$ and $5.0$ correspond to the likelihood ratio thresholds $4.0$ and $7.0$ (p-values $0.0455$ and $0.008$).
As Figure 7 shows, we achieve, on average, an AIC reduction of $28.0$, with improvements reaching as high as $121.8$ in some categories. These results confirm that our approach performs consistently well across a wide range of categories, which span both durable and perishable goods.
Figure 7: Fit improvements relative to plain logit across 40 product categories. For each category, we compute $\Delta AIC$ as the $AIC$ of the lowest-$AIC$ specification that includes texts or images minus that of the plain logit model.
Category Data Type Model Type $\Delta$ AIC $\Delta$ HHI
1. Clothing Active Descriptions ST -30.4 129.0%
2. Clothing Shirts Reviews USE -19.1 116.7%
3. Clothing Underwear Images Resnet50 -18.5 88.2%
4. Clothing Sleep Reviews ST -18.5 93.6%
5. Clothing Tops & Blouses Descriptions USE -18.5 94.6%
6. Electronics Cables Titles USE -18.4 240.0%
7. Electronics Accessories Reviews TFIDF -19.8 10.7%
8. Electronics Keyboards Images VGG19 -20.0 88.2%
9. Electronics Memory Cards Images VGG16 -109.7 103.5%
10. Electronics Tablets Images Resnet50 -121.8 120.9%
11. Electronics Monitors Reviews USE -31.6 197.3%
12. Electronics Headphones Descriptions COUNT -26.0 1.2%
13. Electronics Media Players Reviews TFIDF -66.9 42.7%
14. Groceries Water Descriptions COUNT -22.5 123.0%
15. Groceries Coffee Images VGG16 -17.2 45.6%
16. Groceries Tea Titles ST -20.0 145.2%
17. Groceries Chips Reviews ST -13.5 61.9%
18. Household Aromatherapy Titles COUNT -19.7 190.7%
19. Household Batteries Reviews TFIDF -20.1 4.8%
20. Household Trash Bags Titles ST -24.6 93.5%
21. Household Paper Towels Titles TFIDF -28.9 188.4%
22. Home Sheets & Pillowcases Descriptions USE -22.3 121.2%
23. Bedroom Beds Reviews USE -17.1 69.4%
24. Bedroom Mattresses Titles USE -17.4 66.3%
25. Kitchen Food Storage Images Resnet50 -27.7 133.2%
26. Office Folders Titles USE -29.7 199.7%
27. Office Paper Images Resnet50 -19.4 11.4%
28. Office Markers Images VGG19 -22.3 228.7%
29. Office Pens Titles ST -14.4 241.0%
30. Office Printer Supplies Titles TFIDF -23.2 247.7%
31. Pet Cat Food Titles ST -16.9 130.6%
32. Pet Cat Litter Titles ST -21.9 126.9%
33. Pet Cat Snacks Images Xception -19.5 142.3%
34. Pet Dog Food Titles COUNT -19.2 65.6%
35. Pet Dog Treats Reviews ST -22.9 163.9%
36. Game Consoles Nintendo Titles COUNT -13.1 118.3%
37. Video Games Nintendo Images Resnet50 -36.7 26.2%
38. Video Games PC Images VGG19 -18.6 43.9%
39. Video Games PS4 Titles TFIDF -32.1 159.6%
40. Video Games Xbox Reviews USE -40.5 120.6%
Table 2: Estimation results across 40 categories in Comscore data. For each category, the table shows the selected model and data type yielding the lowest $AIC$, the corresponding $AIC$ improvement relative to the plain logit model, and the relative change in the $HHI$ of the predicted diversion ratios relative to the plain logit model.
4.2.2 Implications for Substitution
Next, we examine how incorporating unstructured data into demand estimation affects our understanding of product substitution. While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how diversion ratios estimated by our approach deviate from the restrictive patterns imposed by the plain logit model.
The assumption of independent utilities in the plain logit model implies that diversion ratios depend only on market shares and not on product similarities. Consequently, this model often fails to identify close substitutes, leading to overly flat diversion ratios (Conlon et al., 2023). By contrast, if our approach better identifies close substitutes, it should produce more unequal diversion ratios. To quantify this, we compute diversion ratios using the definition from Section 3 and calculate the Herfindahl-Hirschman Index ($HHI$) of predicted diversion ratios. [^23] Specifically, we compute the $HHI$ across $(J-1)$ conditional second choice probabilities for each first choice and then average the resulting $HHI$ measure across all $J$ possible first choices. We expect our approach to yield a higher $HHI$ than the plain logit model.
The last column in Table 2 presents the results. [^24] In some categories, our estimates of the price coefficient $\alpha$ are positive, likely due to correlation between prices and unobserved demand shocks, even after accounting for product fixed effects. While instrumental variables could address this, we do not pursue this approach, as addressing price endogeneity across many product categories is orthogonal to the contribution of our paper. Instead, we focus on counterfactuals, such as the diversion ratios from removing a product, which remain well-defined even when the price coefficient is positive. Across categories, our approach more than doubles the $HHI$ of diversion ratios, with an average increase of 115%. Additionally, $HHI$ rises in all 40 product categories, with some increases approaching 250%. These findings confirm that our approach produces significantly more variable diversion ratios across the board.
4.2.3 Relevance of Different Data Types

### v3

@@ 3.3 Validation Using Second-Choice Data
The Inmate (M) 0.325 The Inmate (M) 0.153 The Inmate (M) 0.162 The Inmate (M) 0.173
The Housemaid (M) 0.250 Don’t Believe (S) 0.149 The Housemaid (M) 0.151 The Housemaid (M) 0.169
Don’t Believe (S) 0.088 The Housemaid (M) 0.145 Don’t Believe (S) 0.136 Don’t Believe (S) 0.115
Art of Letting Go (S) 0.066 Dopamine Detox (S) 0.137 Dopamine Detox (S) 0.115 Court of Ravens (F) 0.107
Serpent & Wings (F) 0.066 Art of Letting Go (S) 0.110 Art of Letting Go (S) 0.106 Serpent & Wings (F) 0.107
Dopamine Detox (S) 0.063 Court of Ravens (F) 0.100 Court of Ravens (F) 0.102 Dopamine Detox (S) 0.101
Court of Ravens (F) 0.058 Serpent & Wings (F) 0.092 Serpent & Wings (F) 0.100 Art of Letting Go (S) 0.095
The Ritual (M) 0.056 The Ritual (M) 0.059 The Ritual (M) 0.064 The Ritual (M) 0.068
Ashes & Star (F) 0.029 Ashes & Star (F) 0.056 Ashes & Star (F) 0.064 Ashes & Star (F) 0.063
@@
Serpent & Wings (F) 0.275 Please Tell Me (M) 0.159 Please Tell Me (M) 0.184 Please Tell Me (M) 0.176
Court of Ravens (F) 0.243 The Inmate (M) 0.136 The Inmate (M) 0.158 The Housemaid (M) 0.152
The Inmate (M) 0.105 Don’t Believe (S) 0.133 The Housemaid (M) 0.138 The Inmate (M) 0.150
The Ritual (M) 0.082 The Housemaid (M) 0.129 Serpent & Wings (F) 0.123 Court of Ravens (F) 0.117
Please Tell Me (M) 0.080 Dopamine Detox (S) 0.122 The Ritual (M) 0.097 Serpent & Wings (F) 0.104
The Housemaid (M) 0.070 Art of Letting Go (S) 0.098 Court of Ravens (F) 0.086 Don’t Believe (S) 0.086
Don’t Believe (S) 0.052 Court of Ravens (F) 0.089 Don’t Believe (S) 0.081 Dopamine Detox (S) 0.080
Art of Letting Go (S) 0.048 Serpent & Wings (F) 0.082 Art of Letting Go (S) 0.071 Art of Letting Go (S) 0.072
Dopamine Detox (S) 0.045 The Ritual (M) 0.053 Dopamine Detox (S) 0.061 The Ritual (M) 0.063
@@
Implied Substitution Patterns
To illustrate how our approach improves predictions of substitution patterns, in Table 2 we compare predicted second-choice probabilities with their counterparts observed in the experimental data for three of the books.
Panel A examines substitution patterns for the self-help book Dopamine Detox. The two other self-help books are, by far, its closest substitutes in the data. The plain logit model misidentifies these substitutes, incorrectly predicting that people would switch to books Please Tell Me and The Inmate—two popular books with the largest market shares. The attribute-based mixed logit correctly identifies Don’t Believe Everything You Think as the closest substitute but mispredicts the second one, likely due to its over-reliance on estimated fixed effects. By contrast, our review-based model is the only one that correctly predicts all five closest substitutes in the correct order.
Panel B shows a similar example with the substitutes for the mystery book Please Tell Me. The plain logit model mispredicts second-choice probabilities, incorrectly suggesting that the second-closest substitute is a self-help book. By contrast, both the attribute-based mixed logit and our review-based logit capture strong within-genre substitution, correctly identifying the top three closest substitutes. Additionally, the review-based model recognizes that The Inmate and The Housemaid have significantly higher second-choice probabilities than the third-closest substitute—an insight that the attribute-based model misses.
@@
Relation between in-sample and counterfactual performance
Although we select models based on first-choice $AIC$, our experimental data allows us to verify whether this selection algorithm indeed chooses specifications with the best counterfactual performance. Two findings support our choice of $AIC$ as a model selection criterion. First, across all specifications with principal components considered by our model selection algorithm, the correlation between first-choice $AIC$ and counterfactual second-choice $RMSE$ is 0.78. Further, $AIC$ selects the specification that has the lowest second-choice $RMSE$ across all considered specifications. Using $BIC$ instead of $AIC$ selects the same specification, as does a five-fold cross-validation procedure on the first choices. This reassures us that the key results are robust to the choice of model selection method.
@@ 3.4 Implications for Pricing: Merger Simulations
Figure 5: Results of hypothetical merger simulations. For each simulated merger of Dopamine Detox with one other book (horizontal axis), the figure shows the average price increase among the two merging books. The horizontal dashed line represents a hypothetical policy where the decision-maker challenges all mergers expected to raise prices by at least 5%. Appendix Table A2 reports the exact price increase estimates used to construct this graph.
For each pair of books, we compute their prices under two scenarios: (a) when the books are owned by separate publishers competing in a Bertrand-Nash equilibrium, and (b) when both books are owned by the same publisher setting prices by solving a joint first-order condition. In both cases, publishers take the prices of the other eight books as given, fixed at \$5, and have zero marginal costs. [^19] We fix the prices of the other books at \$5, the average value in the experimental dataset. We do not optimize prices of other books because the model does not include an outside option. Consumer choices remain unchanged if all prices increase by the same amount, thus leading to multiple equilibria.
We recognize that limiting the choice set to ten books makes our analysis somewhat artificial, as major publishers typically manage vast assortments with hundreds of thousands of titles. Nevertheless, our simulations offer a natural way to evaluate how estimated substitution patterns translate into pricing decisions. This application is also policy-relevant given the high-profile mergers among major book publishers in the past few decades, which have drawn attention and regulatory scrutiny. [^20] In 2022, a federal judge blocked the proposed merger between Penguin Random House and Simon & Schuster, citing concerns that it would stifle competition (U.S. Department of Justice, 2022). This merger would have reduced the “Big Five” publishers to the “Big Four” and significantly increased the concentration of the publishing market. Prior to that, Penguin Books and Random House merged in 2013 to form Penguin Random House—the largest trade book publisher globally.
@@
As discussed above, the plain logit model fails to capture strong within-genre substitution. Consequently, it incorrectly identifies as the closest substitutes for Dopamine Detox the three mystery books, rather than other self-help ones. This misclassification affects the predicted price changes in merger simulations. Specifically, the plain logit model overstates the price increase when Dopamine Detox is merged with the mystery book Please Tell Me or The Housemaid, while understating the price increase when merged with its true closest substitutes, The Art of Letting Go or Don’t Believe Everything You Think. These discrepancies are substantial: for within-genre mergers, the plain logit predicts modest price increases of 3% and 3.7%, whereas our review-based model estimates them to be 5.6% and 8.9%—approximately 2 to 2.5 times higher (see Appendix Table A2).
The attribute-based mixed logit model produces predictions more aligned with the review-based model. In most cases, both models deviate from the plain logit in the same direction, yielding similar or even identical predicted price increases. However, notable discrepancies remain—for instance, in the case of the self-help books The Art of Letting Go or Don’t Believe Everything You Think, the attribute-based model underestimates the price increase by approximately 25% compared to the review-based model.
@@ 4 Application to Online Retail Data
Next, we apply the same approach to choice data from several online markets. Our first goal is to show that this approach can be applied broadly across categories without being tailored. Our second goal is to identify which text and image data best predict substitution patterns in various categories, thus offering practical guidance on what types of unstructured data researchers should collect for demand estimation.
@@ 4.1 Data
We use purchase data from the 2019-2020 Comscore Web Behavior Panel. Specifically, we use the dataset constructed by Greminger et al. (2023), who classify over 12 million unique products from Amazon.com—browsed or purchased by Comscore panelists—into narrowly defined categories. This dataset also matches purchases with daily product price histories obtained from the third-party database Keepa.com. [^21] We do not observe product rankings, so we do not include them in our demand models. We focus on Amazon.com due to the large volume of Amazon purchases in the Comscore dataset.
@@
We combine purchase data with image and text data we collected from Amazon product detail pages (Figure 6). We use the default product photo for image embeddings. For text embeddings, we use product titles, product descriptions (i.e., bullet points describing the product), and 100 most recent reviews for each product. In estimation, we treat text and image embeddings as fixed over time. While sellers could theoretically modify images and textual descriptions to reposition their products in response to unobserved demand shocks, we find that product images, titles, and descriptions very rarely change over time (see Appendix F).
We apply our method to 40 product categories spanning clothing (shirts, blouses, underwear, sleepwear), household goods (paper towels, trash bags, batteries), office supplies (pens, markers, printing paper), groceries (tea, coffee, bottled water), pet food (wet food, treats), electronics (tablets, monitors, headphones, memory cards, media players), and video games (PC, Nintendo, Xbox, PlayStation). We select the 15 most-purchased products in each category and retain only those categories where these products collectively account for at least 2,000 purchases. This criterion ensures that we observe enough purchases to estimate both product-fixed effects and substitution patterns. [^22] We omit products purchased fewer than 10 times and those with missing price data, which means some categories in our sample include fewer than 15 products. The average product price in selected categories is \$43.
For four electronics categories, we collect standard attributes from product detail pages to compare our approach with attribute-based mixed logit. Specifically, we collect 18 attributes for “Tablets,” 13 attributes for “Monitors,” 20 attributes for “Memory Cards,” and 14 attributes for “Headphones” (see Appendix E for a full list).
4.2 Estimation Results
We apply our approach to each product category and compare it to the plain logit model. In the four electronics categories for which we collected attribute data, we also compare it to the attribute-based mixed logit model. Since we have a relatively large number of attributes in these categories, we reduce their dimensionality via PCA and apply Algorithm 1 for model selection. [^23] We normalize each attribute to have mean zero and variance one before applying PCA.
[…]
Figure 7: Estimated diversion ratios to closest substitutes. This figure plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$ in each category.
Appendix Table A4 shows results from the electronics categories with attribute data. In all four categories, unstructured data meaningfully improves model fit relative to observed attributes, suggesting that our approach captures information about substitution patterns beyond that contained in standard attributes. This finding is particularly noteworthy given that one might expect technical specifications of electronics products—captured by our extensive list of attributes—to be highly relevant for consumer choices.
Next, we examine the estimated substitution patterns. While we do not observe the “ground truth” diversion ratios as we did in the experiment, we can still assess how estimated diversion ratios deviate from those in the plain logit model. A well-known limitation of plain logit is that diversion ratios depend only on market shares and not on attribute similarities, so this model often fails to identify close substitutes, producing overly flat diversions (Conlon et al., 2023). If our approach recovers substitution better, it should produce more variable diversion ratios.
Figure 7 plots the estimated diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$. Consistent with the intuition above, the plain logit yields relatively small diversion ratios, about 22% on average, while our approach increases this average substantially to 47%. In some categories, our approach estimates diversion ratios as high as 60-80%, which the plain logit does not produce. These results confirm that our method generates significantly more variable diversion ratios, suggesting it better identifies which products are close substitutes.
Apart from showing greater variability, the diversion ratios predicted by our method are also more intuitive. For example, Appendix Tables A5 and A6 report predicted diversions for tablets. Because the plain logit relies on product fixed effects, it predicts that if any tablet is removed, consumers substitute to the most popular alternatives (e.g., Fire 7 or Fire HD10), producing unrealistically flat diversions. By contrast, the Description ST model chosen by our model selection algorithm yields more intuitive diversions: the two kids’ tablets are estimated to be close substitutes (Fire Kids and Fire HD8 Kids), the two iPads are close substitutes (iPad 9.7 and iPad 10.2), and so on.
4.3 Relevance of Different Data Types